# Decoding Noise in Nanofluidic Systems: Adsorption versus Diffusion Signatures in Power Spectra

### v1

@@ Appendix C Computational methods
Each step is carried out as follows. For each particle, we first check if it is already adsorbed. If it is, a random number is sampled from a uniform distribution $r_{1}\sim U(0,1)$. If $r_{1}<p_{\rm off}$, the particle unadsorbs, and if $r_{1}>p_{\rm off}$, it remains adsorbed. If the particle was not previously adsorbed, but is in the channel region, a second random number $r_{2}\sim U(0,1)$ is generated. If $r_{2}<p_{\rm on}$, the particle adsorbs, and if $r_{2}>p_{\rm on}$, it remains unadsorbed. If it remains unadsorbed, or if it was initially in the reservoirs, the particle moves according to:
@@ D.1 Diffusion Slower than Adsorption/Desorption
Figure 8: Slow diffusion, fast adsorption/desorption – Case 1 (continued): $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=60/\tau_{0}$, and $k_{\rm off}=20/\tau_{0}$. Note that the main difference compared to Fig. 5 is that relative ordering of the rates of adsorption and desorption are now reversed, here $k_{\rm on}\gg k_{\rm off}$. Labels are similar to Fig. 5.
Now, when the adsorption rate is faster than desorption and binding is favored, the total PSD lies further from the purely diffusive line (dotted) at high frequencies. In this region the free PSD lies closer to the pure adsorption line (dash-dot). These changes can be reconciled with the fact that increasing the adsorption rate makes binding more favorable, so particles adsorb more quickly on entering the channel, on average. This suppresses the high frequency $-3/2$ scaling due to the diffusive motion of particles across boundaries in the total PSD, and also increases the importance of the adsorption process to the free PSD, causing its scaling to tend to $-2$.
Figure 9: All physical processes have comparable timescales – Case 3: $\tau_{\rm cha}\sim\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$ and $k_{\rm on}=k_{\rm off}=0.4/\tau_{0}$.
@@ D.2 Comparable Timescales
Spectra for the case where diffusion, adsorption, and desorption all have similar timescales is shown in Fig. 9. Here, the two corner frequencies and intermediate scaling we could observe clearly in the previous two cases collapse, with all three PSDs showing only one plateau and corner frequency. This illustrates how a complex physical situation with overlapping timescales could be difficult to interpret experimentally. While there are four timescales and two distinct processes (adsorption and diffusion) in the system, they are masked by the simple shape of the PSD.
However, our model and simulations provide a possible means of distinguishing this case from the similar-looking cases of pure diffusion and pure adsorption. The total PSD is close to the pure diffusive case but at intermediate frequencies transiently follows the adsorption case with a scaling of $-2$. The bound PSD, while similar to the case of pure adsorption, has a higher plateau. The free PSD resembles the case of pure diffusion, but its slope is steeper than $-3/2$ at high frequencies as it is still affected by adsorption.
@@
Figure 10: Adsorption much faster than desorption – Case 4: $\tau_{\rm des}\gg\tau_{\rm sys}\sim\tau_{\rm cha}\gg\tau_{\rm ads}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=4/\tau_{0}$, and $k_{\rm off}=0.04/\tau_{0}$.
@@ D.3 Separation of Adsorption and Desorption Timescales
Representative results are shown in Fig. 10. The PSDs now have more complex shapes. The total PSD shows at several scaling regimes, with a corner at the diffusive timescale and a kink at the adsorption timescale. Above the adsorption timescale, it scales as approximately $-3/2$. The bound PSD also shows at least two separate scalings and corner frequencies. At high frequencies, it resembles the pure adsorption case with a $-2$ scaling. The free PSD is very close to the case of pure adsorption (telegraph noise) but its slope at high frequencies is less steep than $-2$. Interestingly, the diffusive timescale along the channel is the only timescale not corresponding to a change in slope for these PSDs - this may be because very few (if any) particles are able to diffuse along the channel without adsorption.

### v2

@@ Appendix C Computational methods
Each step is carried out as follows. For each particle, we first check if it is already adsorbed. If it is, a random number is sampled from a uniform distribution $r_{1}\sim U(0,1)$. If $r_{1}<p_{\rm off}$, the particle un-adsorbs, and if $r_{1}>p_{\rm off}$, it remains adsorbed. If the particle was not previously adsorbed, but is in the channel region, a second random number $r_{2}\sim U(0,1)$ is generated. If $r_{2}<p_{\rm on}$, the particle adsorbs, and if $r_{2}>p_{\rm on}$, it remains unadsorbed. If it remains unadsorbed, or if it was initially in the reservoirs, the particle moves according to:
@@
Figure 9: Slow diffusion, fast adsorption/desorption – Case 1 (continued): $\tau_{\rm cha}\gg\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=375/\tau_{0}$, and $k_{\rm off}=125/\tau_{0}$. Note that the main difference compared to Fig. 5 is that relative ordering of the rates of adsorption and desorption are now reversed, here $k_{\rm on}\gg k_{\rm off}$. Labels are similar to Fig. 5.
@@
In the 3D simulations, the procedure is exactly as above, except each particle has three position coordinates and diffuses in a cubic box of side length $L$. The “channel region”, or sensing region, is a cubic box of length $L_{0}$ at the center.
@@ Appendix D Additional results in the full case
Figure 10: All physical processes have comparable timescales – Case 3: $\tau_{\rm cha}\sim\tau_{\rm ads}\sim\tau_{\rm des}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$ and $k_{\rm on}=k_{\rm off}=2.5/\tau_{0}$.
@@ D.1 Diffusion Slower than Adsorption/Desorption
Now, when the adsorption rate is faster than desorption and binding is favored, the total PSD lies further from the purely diffusive line (dotted) at high frequencies. In this region the free PSD lies closer to the pure adsorption line (dash-dot). These changes can be reconciled with the fact that increasing the adsorption rate makes binding more favorable, so particles adsorb more quickly on entering the channel, on average. This suppresses the high frequency $f^{-3/2}$ scaling due to the diffusive motion of particles across boundaries in the total PSD, and also increases the importance of the adsorption process to the free PSD. Nonetheless, at very high frequencies both toal and free PSDs eventually recover their $f^{-3/2}$ scaling.
@@ D.2 Comparable Timescales
Spectra for the case where diffusion, adsorption, and desorption all have similar timescales are shown in Fig. 10. Here, the two corner frequencies and intermediate scaling we could observe clearly in the previous two cases collapse, with all three PSDs showing only one plateau and corner frequency. This illustrates how a complex physical situation with overlapping timescales could be difficult to interpret experimentally. While there are four timescales and two distinct processes (adsorption and diffusion) in the system, they are masked by the simple shape of the PSD.
However, our model and simulations provide a possible means of distinguishing this case from the similar-looking cases of pure diffusion and pure adsorption. The total PSD is close to the pure diffusive case but at intermediate frequencies transiently follows the adsorption case with a scaling of $f^{-2}$. The bound PSD, while similar to the case of pure adsorption, has a higher plateau. The free PSD resembles the case of pure diffusion, but its slope is steeper than $f^{-3/2}$ at intermediate frequencies as it is still affected by adsorption (until it recovers the $f^{-3/2}$ scaling at high frequencies.
@@
Figure 11: Adsorption much faster than desorption – Case 4: $\tau_{\rm des}\gg\tau_{\rm sys}\sim\tau_{\rm cha}\gg\tau_{\rm ads}$. Left: total PSD. Middle: bound PSD. Right: free PSD. We use $L=2\ell_{0}$, $k_{\rm on}=25/\tau_{0}$, and $k_{\rm off}=0.25/\tau_{0}$.
@@ D.3 Separation of Adsorption and Desorption Timescales
Representative results are shown in Fig. 11. The PSDs now have more complex shapes. The total PSD shows several scaling regimes, with a corner at the diffusive timescale and a kink at the adsorption timescale. Above the adsorption timescale, it scales as $f^{-3/2}$. The bound PSD also shows at least two separate scalings and corner frequencies. At high frequencies, it resembles the pure adsorption case with a $f^{-2}$ scaling. The free PSD is very close to the case of pure adsorption (telegraph noise) but its slope at high frequencies eventually returns to $f^{-3/2}$. Interestingly, the diffusive timescale along the channel is the only timescale not corresponding to a change in slope for these PSDs - this may be because very few (if any) particles are able to diffuse along the channel without adsorption.